OpenStudy (anonymous):

Make this a perfect square? w^2+10w+ blank

OpenStudy (unklerhaukus):

(w+a)^2 = (w+a)*(w+a) = w^2+aw+wa+a^2 = w^2+2aw+a^2

OpenStudy (anonymous):

im so sorry but im confused

OpenStudy (unklerhaukus):

well if you compare w^2+2aw+a^2 & w^2+10w+__ can you find a?

OpenStudy (anonymous):

would it be 5...?

OpenStudy (unklerhaukus):

yeah 2a =10 so a = 5

OpenStudy (unklerhaukus):

now you have to square this to get the blank

OpenStudy (anonymous):

so w^2 +10w +25?

OpenStudy (unklerhaukus):

√√√

OpenStudy (anonymous):

okay thanks!(: <3